find three consecutive positive odd integers such that the product of the two smaller exceeds the largest by 26
Answer by Alan3354(69443)

find three consecutive positive odd integers such that the product of the two smaller exceeds the largest by 26
-------------
Use n-1, n+1 and n+3
n^2-1 = n+3 + 26
n^2 - n - 30 = 0
etc